# The Cramér-Rao approach and global quantum estimation of bosonic states

### v4

@@ Appendix D Proof of Theorem 1
$R(M_{\rho^{\prime}})=\int_{\mathop{\rm SU}(2)}\mathrm{Tr}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|U_{1/2,g}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|U_{1/2,g}^{\dagger}\cdot(2j+1)\mathrm{Tr}\rho U_{j,g}\rho^{\prime}U_{j,g}^{\dagger}\nu(dg)$ (115)
$=$ $\int_{\mathop{\rm SU}(2)}(2j+1)\mathrm{Tr}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho(U_{1/2,g}\otimes U_{j,g})|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime}(U_{1/2,g}\otimes U_{j,g})^{\dagger}\nu(dg)$ (116)
$=$ $(2j+1)\mathrm{Tr}(P_{j+\frac{1}{2}}+P_{j-\frac{1}{2}})(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho)(P_{j+\frac{1}{2}}+P_{j-\frac{1}{2}})\int_{\mathop{\rm SU}(2)}(U_{1/2,g}\otimes U_{j,g})(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime})(U_{1/2,g}\otimes U_{j,g})^{\dagger}\nu(dg).$ (117)
@@
$R(M_{\rho^{\prime}})=$ $\frac{2j+1}{2j+2}\mathrm{Tr}P_{j+\frac{1}{2}}(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho)(\mathrm{Tr}P_{j+\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime})$ (120)
$+\frac{2j+1}{2j}\mathrm{Tr}P_{j-\frac{1}{2}}(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho)(\mathrm{Tr}P_{j-\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime})$ (121)
@@
Appendix E The error function for the global estimate on the unitary model with the geometric distribution
@@
$1-R(M_{n})=1-\frac{n+1}{n+2}\Big{(}\frac{1}{n+1}+\frac{1}{n+1}(\frac{nr^{n+1}+1}{r^{n+1}-1}-\frac{1}{r-1})\Big{)}=1-\Big{(}\frac{1}{n+2}+\frac{1}{n+2}(\frac{nr^{n+1}+1}{r^{n+1}-1}-\frac{1}{r-1})\Big{)}$ (123)
$=$ $1-\Big{(}\frac{1}{n+2}+\frac{1}{n+2}(\frac{n+r^{-n-1}}{1-r^{-n-1}}-\frac{1}{r-1})\Big{)}$ (124)
$=$ $1-\Big{(}\frac{1}{n+2}+\frac{1}{n+2}(n+O(r^{-n-1})-\frac{1}{r-1})\Big{)}$ (125)
$=$ $\frac{1}{n+2}\Big{(}n+2-1-(n+O(r^{-n-1})-\frac{1}{r-1})\Big{)}$ (126)
$=$ $\frac{1}{n+2}\Big{(}1-\frac{1}{r-1}+O(r^{-n-1})\Big{)}$ (127)
$=$ $\frac{r}{(n+2)(r-1)}+O(r^{-n-1})$ (128)
$=$ $\frac{r}{n(r-1)}-\frac{2}{n^{2}(r-1)}+O(\frac{1}{n^{3}})+O(r^{-n-1}).$ (129)

### v5

@@ Appendix E Proof of Theorem 1
$R(M_{\rho^{\prime}})=\int_{\mathop{\rm SU}(2)}\mathrm{Tr}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|U_{1/2,g}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|U_{1/2,g}^{\dagger}\cdot(2j+1)\mathrm{Tr}\rho U_{j,g}\rho^{\prime}U_{j,g}^{\dagger}\nu(dg)$
$=$ $\int_{\mathop{\rm SU}(2)}(2j+1)\mathrm{Tr}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho(U_{1/2,g}\otimes U_{j,g})|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime}(U_{1/2,g}\otimes U_{j,g})^{\dagger}\nu(dg)$
$=$ $(2j+1)\mathrm{Tr}(P_{j+\frac{1}{2}}+P_{j-\frac{1}{2}})(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho)(P_{j+\frac{1}{2}}+P_{j-\frac{1}{2}})$
×$\quad\times\int_{\mathop{\rm SU}(2)}(U_{1/2,g}\otimes U_{j,g})(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime})(U_{1/2,g}\otimes U_{j,g})^{\dagger}\nu(dg).$ (120)
@@
$R(M_{\rho^{\prime}})=$ $\frac{2j+1}{2j+2}\mathrm{Tr}P_{j+\frac{1}{2}}(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho)(\mathrm{Tr}P_{j+\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime})$
$+\frac{2j+1}{2j}\mathrm{Tr}P_{j-\frac{1}{2}}(|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho)(\mathrm{Tr}P_{j-\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho^{\prime})$
@@
Appendix F The error function for the global estimate on the unitary model with the geometric distribution
@@
$1-R(M_{n})=1-\frac{n+1}{n+2}\Big{(}\frac{1}{n+1}+\frac{1}{n+1}(\frac{nr^{n+1}+1}{r^{n+1}-1}-\frac{1}{r-1})\Big{)}=1-\Big{(}\frac{1}{n+2}+\frac{1}{n+2}(\frac{nr^{n+1}+1}{r^{n+1}-1}-\frac{1}{r-1})\Big{)}$
$=$ $1-\Big{(}\frac{1}{n+2}+\frac{1}{n+2}(\frac{n+r^{-n-1}}{1-r^{-n-1}}-\frac{1}{r-1})\Big{)}=1-\Big{(}\frac{1}{n+2}+\frac{1}{n+2}(n+O(r^{-n-1})-\frac{1}{r-1})\Big{)}$
$=$ $\frac{1}{n+2}\Big{(}n+2-1-(n+O(r^{-n-1})-\frac{1}{r-1})\Big{)}=\frac{1}{n+2}\Big{(}1-\frac{1}{r-1}+O(r^{-n-1})\Big{)}$
$=$ $\frac{r}{(n+2)(r-1)}+O(r^{-n-1})=\frac{r}{n(r-1)}-\frac{2}{n^{2}(r-1)}+O(\frac{1}{n^{3}})+O(r^{-n-1}).$ (124)